OpenStudy (anonymous):

Using the substitution method to solve the following system of equations. 4x – y – 3z = –3 x – 2y – 2z = –3 x + y + 3z = –2

OpenStudy (anonymous):

ok subtract equation 3 from 2 x-2y-2z-(x+y+3z)=-3-(-2) -3y-5z=-1 ...........A multiply the equation 2 by 2 4x-4y-8z=-6 subtract this from equation 1 4x-y-3z-(4x-4y-8z)=-3-(-6) 3y+5z=3 ..............B solve equations A and B

OpenStudy (anonymous):

@sammyrayepatton can you solve equation A and B.

OpenStudy (anonymous):

@sammyrayepatton waiting for your reply !

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

you 're welcome :)